Opened 9 years ago

Closed 8 years ago

Last modified 2 years ago

#277 closed flight processing (fixed)

GB08/02, flight day 281/2009, Delamere

Reported by: mark1 Owned by: knpa
Priority: alpha 4 high Milestone: 2009 Data processing completion
Component: Archiving Keywords:
Cc: Other processors: mark1

Description (last modified by knpa)

Data location: ~arsf/arsf_data/2009/flight_data/uk/GB08_02-2009_281_Delamere

Data arrived from ARSF via network transfer on 9th October 2009.

Scientific objective: Integrating Airborne Lidar and Terrestrial Laser Scanning to characterise phenological changes in forest vegetation

Priority: alpha-4-High

PI: M Danson

Sensors:

  • Eagle (12/04/2010)
  • Hawk (12/04/2010)
  • Leica LIDAR (24/11/2009)

Change History (25)

comment:1 Changed 9 years ago by mark1

Have renamed Hawk281* to SWIR281*

comment:2 Changed 9 years ago by benj

Line 1 should probably be ignored - labelled VOID on the logsheet, second line is a repeat.

Hawk nav files appear to have no sync at all on any of them. Processing complains about missing sync (line 3 segfaults with my quick test, need to work out why), running getsync.py -t *.nav on the Hawk directory suggests all of the nav files are partially corrupt and none have sync.

comment:3 Changed 9 years ago by knpa

  • Owner set to knpa
  • Status changed from new to assigned

Navigation processed by emca, starting main body of processing.

comment:4 Changed 9 years ago by mggr

Re: #281 (Hawk frame rates are double what they should be), the following lines are affected and the .hdr file will need to be edited to halve the fps number:

uk/GB08_02-2009_281_Delamere/hawk/SWIR281-09-1.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB08_02-2009_281_Delamere/hawk/SWIR281-09-2.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB08_02-2009_281_Delamere/hawk/SWIR281-09-4.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB08_02-2009_281_Delamere/hawk/SWIR281-09-5.hdr: predicted length half recorded length, frame rate doubling problem?

comment:5 Changed 9 years ago by knpa

LIDAR completed and awaiting delivery check

comment:6 Changed 9 years ago by emca

  • Owner changed from knpa to emca
  • Status changed from assigned to new

comment:7 Changed 9 years ago by emca

LIDAR delivery check complete.
Changes made to Read_me & screenshots.

Data quality report needs to be updated with the latest calibration information before data can be sent.

comment:8 Changed 9 years ago by knpa

  • Owner changed from emca to knpa
  • Status changed from new to accepted

comment:9 Changed 9 years ago by knpa

LIDAR DELIVERED

comment:10 Changed 9 years ago by mark1

  • Other processors set to mark1

Started Hyperspectral processing

comment:11 Changed 9 years ago by mark1

No dark frames on the hawk files.

comment:12 Changed 9 years ago by mark1

Hawk line SWIR281-09-3 has incorrect number of lines in hdr. Changed from 10 to 3230.

comment:13 Changed 9 years ago by mark1

SCT offsets:

Raw Eagle lineSCT Raw Hawk SCT
-2 -0.02 -2 -6.29
-3 0.01 -3 -5.02
-4 0.01 -4 -5.01
-5 0.01 -5 -4.87

comment:14 Changed 9 years ago by mark1

Hyperspectral delivery created at
~airborne/workspace/GB08_02-2009_281_Delamere/delivery/20100409/GB08-02

comment:15 Changed 9 years ago by benj

Delivery checked, made a couple of minor updates to the readme and updated the copyright notice date to 2010. Hawk looks slightly worse than usual to me, but this is likely to be a result of having to substitute dark frames from another flight.

Ready for rsyncing and delivery.

comment:16 Changed 9 years ago by mark1

  • Description modified (diff)

Hyperspectral delivery dispatched to PI on 12/04/2010

comment:17 Changed 9 years ago by mark1

All data delivered and rsynced over to main ~arsf repository.

comment:18 Changed 9 years ago by knpa

  • Component changed from Processing: general to Archiving

comment:19 Changed 8 years ago by knpa

  • Status changed from accepted to assigned

Beginning archiving.

comment:20 Changed 8 years ago by knpa

  • Description modified (diff)

comment:21 Changed 8 years ago by knpa

  • Description modified (diff)

No webcam for this project.

No .reg file for lidar.

comment:22 Changed 8 years ago by knpa

Finished preparing for archiving.

Made available to NEODC.

comment:23 Changed 8 years ago by knpa

Downloaded and backed up NEODC.

Moved to archive disk.

comment:24 Changed 8 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Deleted workspace version.

Closing ticket.

comment:25 Changed 2 years ago by asm

Archiving

Re-processed hyperspectral data has been archived and everythin is ready to be downloaded from CEDA. Will close this ticket.

Note: See TracTickets for help on using tickets.